AP Stats: Probability Exam Solved
Correctly To Score A+ 2025!!!
Probability CORRECT ANSWERS The collection of all possible outcomes of a chance
experiment in the sample space for the experiment
Event CORRECT ANSWERS Any collection of outcomes from the sample space
Simple Event CORRECT ANSWERS An event with exactly one outcome
Venn diagram CORRECT ANSWERS Rectangle + a circle for each event
Circles and rectangle space should add up to probability of 1
Classical Approach CORRECT ANSWERS If a chance experiment has K outcomes, all
equally likely, then each individual outcome has the probability 1/k
P(E)= (number of outcomes resulting in E)/number of outcomes in sample space
Empirical Approach (Law of Large Numbers) CORRECT ANSWERS The empirical
approach will stabilize at the theoretical value if the chance experiment is repeated
many times
This means that any unusual sequence near,the beginning does not affect the
stabilization towards the theoretical value
Relative Frequency Approach CORRECT ANSWERS For an event E, P(E) is the value
approached by the relative frequency occurrence of E in a long series of chance trials
Subjective Approach CORRECT ANSWERS Use of personal judgement and opinions
to determine probability
BR1: For any event E, CORRECT ANSWERS P(E) is between 0 and 1, INCLUSIVE
BR2: If S is the sample space for an experiment, CORRECT ANSWERS P(S)=1
*BR3: If events E and F are disjoint, CORRECT ANSWERS P(E or F)=P(E)+P(F)
BR4: For any event E (complementary), CORRECT ANSWERS P(E)+P(not E)=1
P(not E)= 1-P(E)
P(E)= 1-P(not E)
Correctly To Score A+ 2025!!!
Probability CORRECT ANSWERS The collection of all possible outcomes of a chance
experiment in the sample space for the experiment
Event CORRECT ANSWERS Any collection of outcomes from the sample space
Simple Event CORRECT ANSWERS An event with exactly one outcome
Venn diagram CORRECT ANSWERS Rectangle + a circle for each event
Circles and rectangle space should add up to probability of 1
Classical Approach CORRECT ANSWERS If a chance experiment has K outcomes, all
equally likely, then each individual outcome has the probability 1/k
P(E)= (number of outcomes resulting in E)/number of outcomes in sample space
Empirical Approach (Law of Large Numbers) CORRECT ANSWERS The empirical
approach will stabilize at the theoretical value if the chance experiment is repeated
many times
This means that any unusual sequence near,the beginning does not affect the
stabilization towards the theoretical value
Relative Frequency Approach CORRECT ANSWERS For an event E, P(E) is the value
approached by the relative frequency occurrence of E in a long series of chance trials
Subjective Approach CORRECT ANSWERS Use of personal judgement and opinions
to determine probability
BR1: For any event E, CORRECT ANSWERS P(E) is between 0 and 1, INCLUSIVE
BR2: If S is the sample space for an experiment, CORRECT ANSWERS P(S)=1
*BR3: If events E and F are disjoint, CORRECT ANSWERS P(E or F)=P(E)+P(F)
BR4: For any event E (complementary), CORRECT ANSWERS P(E)+P(not E)=1
P(not E)= 1-P(E)
P(E)= 1-P(not E)